OpenStudy (anonymous):

Solve 3x^ 2 + 11x + 6 = 0

OpenStudy (ashleyisakitty):

Is that 3x^2?

OpenStudy (anonymous):

Solve 3x^2 + 11x + 6 = 0 sorry wrote it wrong

OpenStudy (shamim):

\[x=\frac{ -11\pm \sqrt{11^{2}-4 \times 6 \times 3}}{ 2 \times 3 }\]

OpenStudy (shamim):

did u get it

OpenStudy (anonymous):

\[3x^2+11x+6=0\] U gotta multiply 3x² and 6, mentally . its 18x² -Now u gotta find two numbers that are factors of 18x² yet they add up to 11x. -11x is the middle term btw Those 2 numbers are 9x and 2x... ((EX: 9x*2x=18x² 9x+2x=11x)) - Now fill those 2 numbers in for 11x 3x²+9x+2x+6=0 (undistribute) 3x(x+3)+2(x+3)=0 (grouping) (3x+2)(x+3)=0 FINAL ANSWER= x= -2/3, x=-3

OpenStudy (anonymous):

the explanation for the last part, is: (3x+2)(x+3)=0 3x+2=0 x+3=0 -2 -2 -3 -3 3x=-2 x=-3 /3 /3 x=-2/3